Suppose a sample of size 21 is taken, resulting in a sample mean of 260. Determine the Point Estimate and the Standard Error for a 90% Confidence Interval for this sample. Round all calculations to two (2) decimal places.
Answer #1

Solution :

Point estimate = 260

Standard error = (s /n) = s / 21